Based on the relative frequency histogram below, which of the following statements is supported by the plot?


1
Expert's answer
2021-12-08T04:25:28-0500


Answer: the IQR of the distribution is about 10


"IQR=Q_3-Q_1"


Q3 class: (3/4)th value in cumulative frequency:

class 15-20

"Q_3=L+c\frac{0.75-cf}{f}=15+5\frac{0.75-0.85}{0.15}=15-3.33=11.67"

Q1 class: (1/4)th value in cumulative frequency:

class 5-10

"Q_1=L+c\frac{0.25-cf}{f}=5+5\frac{0.25-0.5}{0.4}=5-3.125=1.875"


"IQR=11.67-1.875=9.795"